# A note on the scatteredness of reflection orders

### v1

@@ Abstract.
In this note, we characterize affine Coxeter systems among all Coxeter systems in terms of the structure of their reflection orders. For an infinite irreducible system $(W,S)$, we show that affineness is equivalently characterized by the scatteredness of all reflection orders, by the existence of a reflection order of type $\omega+\omega^{*}$, and by a finiteness property of intervals determined by dihedral reflection subgroups. Our proof exploits the geometry of projective roots, the isotropic cone and the universal reflection subgroups in an infinite non-affine Coxeter group.
@@ 1. Introduction
The longest element in a finite Coxeter group and its reduced expressions play notable roles in the study of the Coxeter group’s combinatorics, representation theory, Schubert geometry, and Kazhdan-Lusztig theory. To compensate the absence of longest elements in infinite Coxeter groups, Dyer introduced in [5] reflection orders, which are analogous to the reduced expressions of the longest elements in infinite setting. Reflection orders are total orders on the set of positive roots, and have applications in shellings of the Bruhat order, combinatorial formulas for Kazhdan–Lusztig and R-polynomials, and the completion of the weak order.
@@
With these applications in mind, it is natural to study the structure of reflection orders. Many conjectures concerning their initial intervals remain unresolved. To better understand the combinatorics of such orders, we initiated in [16] a study of their possible order types. There we showed that the order type of a reflection order of an affine Weyl group is the sum of the order type of the natural numbers and its dual, and that these can be further classified by certain deformed Dyck words.
@@
In this paper, we continue this line of research by studying the order types of non-affine infinite irreducible Coxeter groups. Our main result is that, for any non-affine, infinite, irreducible Coxeter group, there always exists a reflection order $\prec$ and a subset $E$ of the positive roots such that the restriction of $\prec$ to $E$ is dense as a linear order. As a corollary, an irreducible Coxeter group is of affine type if and only if all its reflection orders are scattered in the sense of [14]. The proof exploits information regarding the distribution of projective roots of the universal Coxeter group of rank three. Furthermore, we give another two alternative characterizations of affine Weyl groups among all Coxeter groups using reflection orders, i.e. (1) a Coxeter system is affine if and only if it admits a reflection order with order type $\omega+\omega^{*}$ and (2) A Coxeter system is affine if and only if, for every dihedral reflection subgroup $U$ and every reflection order $\prec$, there are only finitely many consecutive pairs of positive roots in $\Phi_{U}^{+}$ whose interval under $\prec$ is infinite.
@@
This paper is organized as follow: In Section 2, we review the necessary background on Coxeter groups and root systems, as well as linear orderings, reflection orders, and projective representations of roots. Section 3 studies an explicit reflection order on a rank-3 universal Coxeter system and determines its order type, which serves as a key tool in the proofs of the main results. The main results are presented and proved in Section 4.
@@ 2.1. Coxeter system
Let $S$ be a set. A Coxeter matrix is a function $m:S\times S\rightarrow\mathbb{Z}_{\geq 1}\cup\{\infty\}$ such that for $s,s^{\prime}\in S$, $m(s,s^{\prime})=1$ if and only if $s=s^{\prime}$, and $m(s,s^{\prime})=m(s^{\prime},s)$. A Coxeter system $(W,S)$ consists of a set $S$ and a group $W$ with the following presentation:
$\langle s\in S\mid\,\forall s,s^{\prime}\in S\,\text{and}\,m(s^{\prime},s)\,\text{is finite},(ss^{\prime})^{m(s,s^{\prime})}=e\rangle$
where $m$ is a Coxeter matrix. The group $W$ is referred to as a Coxeter group and the elements of $S$ are called simple reflections. The cardinality of the set $S$ is known as the rank of $(W,S)$. Notably, $m(s,s^{\prime})$ equals the order of $ss^{\prime}$ (denoted by $\mathrm{ord}(ss^{\prime})$) in $W$. A conjugate of some simple reflection, $wsw^{-1},w\in S,s\in S$, is called a reflection. The set of reflections is denoted by $T$. The length $\ell(w)$ of an element $w\in W$ is defined as the minimum number $k$ such that $w=s_{1}s_{2}\cdots s_{k},s_{i}\in S.$ A subgroup $W^{\prime}$ of $W$ that is generated by a subset $T^{\prime}$ of $T$ is called a reflection subgroup. A reflection subgroup is itself a Coxeter group. Denote by $\chi(W^{\prime})$ the set $\{t\in T\cap W^{\prime}\mid\,\ell(t^{\prime}t)>\ell(t),\forall t^{\prime}\in T\cap W^{\prime},t^{\prime}\neq t\}$. Indeed, $(W^{\prime},\chi(W^{\prime}))$ is a Coxeter system. If the set $S$ has cardinality 2 (i.e. $(W,S)$ has rank 2), then $W$ is a dihedral group. A dihedral reflection subgroup of $W$ is a reflection subgroup of $W$ which is also a dihedral subgroup. It can be shown that every dihedral reflection subgroup is contained in a maximal dihedral reflection subgroup. It can also be shown that a reflection subgroup generated by two distinct reflections must be a dihedral reflection subgroup.
@@ 2.2. The root system of a Coxeter system
Let $(W,S)$ be a rank $n$ Coxeter system where $S=\{s_{1},s_{2},\cdots,s_{n}\}$. Let $V$ be a vector space and let $B$ be a symmetric bilinear form $V\times V\rightarrow\mathbb{R}.$ We call a subset $\Pi$ of $V$ a root basis if
(1) the elements of $\Pi=\{\alpha_{s_{1}},\alpha_{s_{2}},\cdots,\alpha_{s_{n}}\}$ are in bijection with the simple reflections;
(2) $B(\alpha_{i},\alpha_{j})=-\cos\frac{1}{m(s_{i},s_{j})}$ if $m_{s_{i},s_{j}}<\infty$;
(3) $B(\alpha_{i},\alpha_{j})\leq-1$ if $m_{s_{i},s_{j}}=\infty$;
(4) the root basis $\Pi$ is positively independent, i.e. if $\sum_{i=1}^{n}k_{i}\alpha_{s_{i}}=0,k_{i}\geq 0$ for all $i\in\{1,2,\cdots,n\}$, then $k_{i}=0$ for all $i=1,2,\cdots,n.$
The vectors $\alpha_{s_{i}}$ in the root basis are called simple roots. The action of $s\in S$ on $V$ is defined by $s(v)=v-2B(v,\alpha_{s})\alpha_{s}.$ Denote by $\operatorname{{\mathrm{cone}}}(\Pi)$ the set
$\{\sum_{i=1}^{n}k_{i}\alpha_{s_{i}},k_{i}\geq 0,\text{for}\,\text{all}\,\,i=1,2,\cdots,n,\text{and}\,\,k_{i}\neq 0\,\text{for some}\,i\}.$
The root system is the set $\Phi:=W\Pi=\{w(\alpha_{s})\mid\,w\in W,\alpha_{s}\in\Pi\}.$ The set of positive roots $\Phi^{+}$ is defined as the set $\Phi\cap\operatorname{{\mathrm{cone}}}(\Pi)$. The set of negative roots $\Phi^{-}$ is defined as $\Phi^{+}$. The set of roots $\Phi$ is the disjoint union of $\Phi^{+}$ and $\Phi^{-}.$
A root system for a Coxeter group can be constructed concretely as follows. Let be $V$ be an $n$-dimensional Euclidean space with a chosen orthonormal basis $\{\alpha_{s_{1}},\alpha_{s_{2}},\cdots,\alpha_{s_{n}}\}$ whose elements are indexed by elements in $S$. Endow $V$ with another bilinear form B(,)B(-,-) as follows. First require
$B(\alpha_{i},\alpha_{j})=-\cos\frac{\pi}{m(s_{i},s_{j})}$
and extend B(,)B(-,-) bilinearly to $V$. The action of $s\in S$ on $V$ is defined by $s(v)=v-2B(v,\alpha_{s})\alpha_{s}.$ Extending this action to $W$ yields the standard reflection representation of $W$, and then one obtains the root system $\Phi=W\Pi$. It can be shown that $B(w(u),w(v))=B(u,v)$ for all $u,v\in V,w\in W$. Under this construction the root basis is linearly independent and its linear span is $V$.
There exists a bijection between the reflections and the positive roots via the map $w(\alpha_{s_{i}})\mapsto ws_{i}w^{-1}$. Let $\alpha\in\Phi^{+}$ be a root. The reflection in $\alpha$ corresponds to the reflection $s_{\alpha}$ under the above mentioned bijection. The root system of a reflection subgroup $W^{\prime}$ is given by $\Phi^{\prime}=\{\alpha\in\Phi\mid\,s_{\alpha}\in W^{\prime}\}$, where the root basis is given by $\{\alpha\in\Phi^{+}\mid\,s_{\alpha}\in\chi(W^{\prime})\}.$
@@ 2.3. The order type of a total order
Let $(A,\prec_{A})$ and $(B,\prec_{B})$ be two totally ordered sets. They are said to be isomorphic if there exists a bijective map $\phi:A\rightarrow B$ such that
$\phi(a)\prec_{B}\phi(b)\,\,\,\text{if}\,\text{and}\,\text{only}\,\text{if}\,\,\,a\prec_{A}b.$
Clearly isomorphisms among total orders give rise to an equivalence relation on the class of all totally ordered sets. The order type is then a representative of an equivalence class. A totally ordered set is said to have the order type $\theta$ if it is in an equivalence class represented by $\theta$. The order type of the natural numbers $\mathbb{N}$ (with the natural order) is denoted by $\omega$. The order type of a totally ordered set with $n$ elements is denoted by $[n]$. The order type of the rational numbers $\mathbb{Q}$ (with the natural order) is denoted by $\eta$. The backward total order of a totally ordered set $(A,\prec_{A})$ is denoted by $(A,\prec_{A}^{*})$. If the order type of $(A,\prec_{A})$ is $\theta$, then the order type of $(A,\prec_{A}^{*})$ is denoted by $\theta^{*}.$
Let $(I,\prec^{\prime})$ be a totally ordered (index) set. For disjoint totally ordered sets $(A_{i},\prec_{A_{i}}),i\in I$, a natural total order $\prec$ can be defined on their union $\cup_{i\in I}A_{i}$ by requiring that $a\prec b$ if and only if $a,b\in A_{i},a\prec_{A_{i}}b$ or $a\in A_{j},b\in A_{k},j\prec^{\prime}k.$ If $(A_{i},\prec_{A_{i}})$ has the order type $\theta_{i}$, then the order type of $(\cup_{i\in I}A_{i},\prec)$ is denoted by $\sum_{i\in I}\theta_{i}$. Suppose that all $(A_{i},\prec_{A_{i}})$ are isomorphic and the order type of them is $\tau_{1}$ and that the order type of $(I,\prec^{\prime})$ is $\tau_{2}$, then the order type of $(\cup_{i\in I}A_{i},\prec)$ is denoted by $\tau_{1}\cdot\tau_{2}$.
A total order $(A,\prec_{A})$ is said to be dense if, for any two elements $a,b\in A$ with $a\prec_{A}b$, there exists an element $c\in A$ such that $a\prec_{A}c\prec_{A}b$. A total order $(A,\prec_{A})$ is called scattered if the restriction of $\prec_{A}$ to any subset $E$ of $A$ is not dense. The standard reference for the notions mentioned in this subsection is [14].
@@ 2.4. Reflection order
A total order $\prec$ on $\Phi^{+}$ is called a reflection order if, for any $\alpha,\beta\in\Phi^{+}$ with $\alpha\prec\beta$ and any $\gamma=a\alpha+b\beta\in\Phi^{+},$ where $a,b\geq 0$, one has $\alpha\prec\gamma\prec\beta.$ Since there exists a bijection between the set of positive roots and the set of reflections, reflection order can be understood as a total order on the set of reflections. Equivalently, a total order $\prec$ on $\Phi^{+}$ is a reflection order if and only if, when restricted to any maximal dihedral reflection subgroup $(W,\{t_{1},t_{2}\})$, it coincides with one of the following orders:
@@
Let $\prec$ be a reflection order of the Coxeter system $(W,S)$ and Let $s\in S$. One can construct another reflection order, called upper $s$-conjugate of $\prec$ and denoted by $\prec^{s}$, as follows: (1) if $\gamma_{1}\prec\gamma_{2}\prec\alpha_{s}$ and $\gamma_{1}\prec\gamma_{2}$, then $\gamma_{1}\prec^{s}\gamma_{2}$; (2) if $\alpha_{s}\prec\gamma_{1}\prec\gamma_{2}$, then $s(\gamma_{1})\prec^{s}s(\gamma_{2})$; (3) if $\gamma_{1}\prec\alpha_{s}\prec\gamma_{2}$, then $\gamma_{1}\prec^{s}\gamma_{2}$; (4) $\alpha_{s}$ is the maximum element under $\prec^{s}$. This notion is also introduced in [5]. See also Section 5.2 in [1].

### v2

@@ Abstract.
In this note, we characterize affine and non-affine Coxeter systems among all Coxeter systems in terms of the structure of their reflection orders. For an infinite irreducible system $(W,S)$, we show that affineness can be characterized in three equivalent ways: by the scatteredness of all reflection orders, by the existence of a reflection order of type $\omega+\omega^{*}$, and by a finiteness property of intervals determined by dihedral reflection subgroups. We also show that non-affineness can be characterized by the existence of order types $(\omega+\omega^{*})[k]$ for arbitrarily large $k$, obtained by restricting any reflection order to a suitable subset. Our proofs exploit the geometry of projective roots, the isotropic cone, and universal reflection subgroups in infinite non-affine Coxeter groups.
@@ 1. Introduction
In the study of the combinatorics of finite Coxeter groups, as well as in representation theory, Schubert geometry, and Kazhdan–Lusztig theory, the longest element and its reduced expressions play a prominent role. To compensate for the absence of longest elements in infinite Coxeter groups, Dyer introduced reflection orders in [4], which are analogous to reduced expressions of the longest elements in the infinite setting. Reflection orders are total orders on the set of positive roots and have a variety of applications, including shellings of the Bruhat order, combinatorial formulas for Kazhdan–Lusztig and $R$-polynomials, and the completion of the weak order.
@@
With these applications in mind, it is natural to study the structure of reflection orders. However, many conjectures concerning their initial intervals remain unresolved. To better understand the combinatorics and global structure of such orders, we initiated a study of their possible order types in [15]. In that work, we showed that, after removing a finite number of roots, the order type of a reflection order of an affine Weyl group is the sum of the order type of the natural numbers and its dual. Moreover, these order types can be further classified using certain deformed Dyck words.
@@
In this paper, we continue this line of research by studying the order types of non-affine, infinite, irreducible Coxeter groups. One of our main results is that, for any such group, there exists a reflection order $\prec$ and a subset $E$ of the positive roots such that the restriction of $\prec$ to $E$ is dense as a linear order. As a corollary, an irreducible Coxeter group is of affine type if and only if all its reflection orders are scattered in the sense of [13]. The proof exploits information regarding the distribution of the projective roots of the universal Coxeter group of rank three.
Furthermore, we give two alternative characterizations of affine Weyl groups among all Coxeter groups using reflection orders: (1) a Coxeter system is affine if and only if it admits a reflection order of type $\omega+\omega^{*}$, and (2) a Coxeter system is affine if and only if, for every dihedral reflection subgroup $U$ and every reflection order $\prec$, there are only finitely many consecutive pairs of positive roots in $\Phi_{U}^{+}$ whose interval under $\prec$ is infinite.
We also give a reflection-order-theoretic characterization of non-affine Weyl groups among all Coxeter groups. Specifically, we show that an infinite, irreducible Coxeter system $(W,S)$ is non-affine if and only if, for arbitrarily large positive integers $k$ and any reflection order $\prec$, there exists a subset $E\subset\Phi^{+}$ such that the restriction $\prec\!\mid_{E}$ has the order type equal to the sum of $k$ copies of $\omega+\omega^{*}$.
@@
This paper is organized as follows. Section 2 provides a review of the necessary background on Coxeter groups and root systems, as well as on linear orderings, reflection orders, and projective representations of roots. Section 3 studies an explicit reflection order on a rank-3 universal Coxeter system and determines its order type, which serves as a key tool in proving the main results. The main results are presented and proved in Section 4. In addition, Section 4 discusses the possible order types of an infinite-rank, locally finite, irreducible Coxeter system and shows that these order types may differ drastically from those in the finite-rank case.
@@ 2.1. Coxeter system
Let $S$ be a set. A Coxeter matrix is a function
$m\colon S\times S\to\mathbb{Z}_{\geq 1}\cup\{\infty\}$
such that for $s,s^{\prime}\in S$, one has $m(s,s^{\prime})=1$ if and only if $s=s^{\prime}$, and $m(s,s^{\prime})=m(s^{\prime},s)$.
A Coxeter system $(W,S)$ consists of a set $S$ and a group $W$ with presentation
$\langle s\in S\mid\forall\,s,s^{\prime}\in S\text{ with }m(s,s^{\prime})<\infty,\ (ss^{\prime})^{m(s,s^{\prime})}=e\rangle,$
where $m$ is a Coxeter matrix. The group $W$ is called a Coxeter group, and the elements of $S$ are called simple reflections. The cardinality of $S$ is called the rank of $(W,S)$. Notably, $m(s,s^{\prime})$ equals the order of $ss^{\prime}$ in $W$, which we denote by $\mathrm{ord}(ss^{\prime})$.
A conjugate $wsw^{-1}$, where $w\in W$ and $s\in S$, is called a reflection. The set of all reflections is denoted by $T$. The length $\ell(w)$ of an element $w\in W$ is defined as the minimum integer $k$ such that
$w=s_{1}s_{2}\cdots s_{k}\quad\text{with }s_{i}\in S.$
A subgroup $W^{\prime}\subseteq W$ generated by a subset $T^{\prime}\subseteq T$ is called a reflection subgroup. Every reflection subgroup is itself a Coxeter group. We define
$\chi(W^{\prime}):=\{\,t\in T\cap W^{\prime}\mid\ell(t^{\prime}t)>\ell(t)\text{ for all }t^{\prime}\in T\cap W^{\prime},\ t^{\prime}\neq t\,\}.$
Then $(W^{\prime},\chi(W^{\prime}))$ is a Coxeter system.
If the set $S$ has cardinality $2$ (that is, if $(W,S)$ has rank $2$), then $W$ is a dihedral group. A dihedral reflection subgroup of $W$ is a reflection subgroup that is itself dihedral. It can be shown that every dihedral reflection subgroup is contained in a maximal dihedral reflection subgroup. Moreover, any reflection subgroup generated by two distinct reflections is necessarily dihedral.
@@ 2.2. The root system of a Coxeter system
Let $(W,S)$ be a Coxeter system of rank $n$, where
$S=\{s_{1},s_{2},\ldots,s_{n}\}.$
Let $V$ be a real vector space equipped with a symmetric bilinear form
$B\colon V\times V\to\mathbb{R}.$
A subset $\Pi\subseteq V$ is called a root basis if the following conditions hold:
  1. (1)
    The elements
    $\Pi=\{\alpha_{s_{1}},\alpha_{s_{2}},\ldots,\alpha_{s_{n}}\}$
    are in bijection with the simple reflections in $S$.
  2. (2)
    $B(\alpha_{s_{i}},\alpha_{s_{j}})=-\cos\!\left(\frac{\pi}{m(s_{i},s_{j})}\right)$ whenever $m(s_{i},s_{j})<\infty$.
  3. (3)
    $B(\alpha_{s_{i}},\alpha_{s_{j}})\leq-1$ whenever $m(s_{i},s_{j})=\infty$.
  4. (4)
    The set $\Pi$ is positively independent; that is, if
    $\sum_{i=1}^{n}k_{i}\alpha_{s_{i}}=0\quad\text{with}\quad k_{i}\geq 0\text{ for all }i,$
    then $k_{i}=0$ for all $i=1,2,\ldots,n$.
The vectors $\alpha_{s_{i}}$ are called the simple roots. For each $s\in S$, define an action on $V$ by
$s(v)=v-2B(v,\alpha_{s})\alpha_{s}.$
Let $\operatorname{{\mathrm{cone}}}(\Pi)$ denote the set
[…]
The root system associated with $(W,S)$ is defined by
$\Phi:=W\Pi=\{w(\alpha_{s})\mid w\in W,\ \alpha_{s}\in\Pi\}.$
The set of positive roots is
$\Phi^{+}:=\Phi\cap\operatorname{{\mathrm{cone}}}(\Pi),$
and the set of negative roots is $\Phi^{-}:=-\Phi^{+}$. Thus $\Phi$ is the disjoint union of $\Phi^{+}$ and $\Phi^{-}$.
A root system for a Coxeter group can be constructed explicitly as follows. Let $V$ be an $n$-dimensional Euclidean space with an orthonormal basis
$\{\alpha_{s_{1}},\alpha_{s_{2}},\ldots,\alpha_{s_{n}}\}$
indexed by the elements of $S$. Define a symmetric bilinear form B(,)B(-,-) on $V$ by setting
$B(\alpha_{s_{i}},\alpha_{s_{j}})=-\cos\!\left(\frac{\pi}{m(s_{i},s_{j})}\right),$
and extending bilinearly to all of $V$. For $s\in S$, define
$s(v)=v-2B(v,\alpha_{s})\alpha_{s},$
and extend this action to all of $W$. This yields the standard reflection representation of $W$, and the associated root system $\Phi=W\Pi$. It can be shown that
$B(w(u),w(v))=B(u,v)\quad\text{for all }u,v\in V\text{ and }w\in W.$
Under this construction, the root basis $\Pi$ is linearly independent and spans $V$.
There is a bijection between the set of reflections in $W$ and the set of positive roots, given by
$w(\alpha_{s_{i}})\longmapsto ws_{i}w^{-1}.$
For $\alpha\in\Phi^{+}$, let $s_{\alpha}$ denote the corresponding reflection. If $W^{\prime}$ is a reflection subgroup of $W$, its root system is
$\Phi^{\prime}=\{\alpha\in\Phi\mid s_{\alpha}\in W^{\prime}\},$
and its root basis is given by
$\{\alpha\in\Phi^{+}\mid s_{\alpha}\in\chi(W^{\prime})\}.$
@@ 2.3. The order type of a total order
Let $(A,\prec_{A})$ and $(B,\prec_{B})$ be two totally ordered sets. They are said to be isomorphic if there exists a bijection
$\phi\colon A\to B$
such that
$\phi(a)\prec_{B}\phi(b)\quad\text{if and only if}\quad a\prec_{A}b$
for all $a,b\in A$. Isomorphism of totally ordered sets defines an equivalence relation on the class of all total orders.
An order type is a representative of such an equivalence class. A totally ordered set is said to have order type $\theta$ if it belongs to the equivalence class represented by $\theta$. The order type of the natural numbers $\mathbb{N}$ with their usual order is denoted by $\omega$. The order type of a totally ordered set with $n$ elements is denoted by $[n]$. The order type of the rational numbers $\mathbb{Q}$ with their usual order is denoted by $\eta$.
Given a totally ordered set $(A,\prec_{A})$, its backward order is denoted by $(A,\prec_{A}^{*})$, where
$a\prec_{A}^{*}b\quad\text{if and only if}\quad b\prec_{A}a.$
If $(A,\prec_{A})$ has the order type $\theta$, then $(A,\prec_{A}^{*})$ has the order type $\theta^{*}$.
Let $(I,\prec^{\prime})$ be a totally ordered index set. For a family of pairwise disjoint totally ordered sets $(A_{i},\prec_{A_{i}})$, $i\in I$, a natural total order $\prec$ can be defined on the union
$\bigcup_{i\in I}A_{i}$
by declaring that
$a\prec b\quad\text{if and only if}\quad\begin{cases}a,b\in A_{i}\text{ and }a\prec_{A_{i}}b,\quad\text{or}\\ a\in A_{j},\ b\in A_{k}\text{ with }j\prec^{\prime}k.\end{cases}$
If $(A_{i},\prec_{A_{i}})$ has order type $\theta_{i}$ for each $i\in I$, then the order type of $\left(\bigcup_{i\in I}A_{i},\prec\right)$ is denoted by
$\sum_{i\in I}\theta_{i}.$
If, moreover, all $(A_{i},\prec_{A_{i}})$ are isomorphic with common order type $\tau_{1}$, and $(I,\prec^{\prime})$ has order type $\tau_{2}$, then the order type of $\left(\bigcup_{i\in I}A_{i},\prec\right)$ is denoted by
$\tau_{1}\cdot\tau_{2}.$
A totally ordered set $(A,\prec_{A})$ is said to be dense if for any $a,b\in A$ with $a\prec_{A}b$, there exists $c\in A$ such that
$a\prec_{A}c\prec_{A}b.$
A totally ordered set $(A,\prec_{A})$ is called scattered if it contains no dense subset; equivalently, the restriction of $\prec_{A}$ to any subset of $A$ is not dense.
The standard reference for the notions introduced in this subsection is [13].
@@ 2.4. Reflection order
A total order $\prec$ on $\Phi^{+}$ is called a reflection order if, for any $\alpha,\beta\in\Phi^{+}$ with $\alpha\prec\beta$ and any $\gamma=a\alpha+b\beta\in\Phi^{+},$ where $a,b\geq 0$, one has $\alpha\prec\gamma\prec\beta.$ Since there exists a bijection between the set of positive roots and the set of reflections, a reflection order can be understood as a total order on the set of reflections. Equivalently, a total order $\prec$ on $\Phi^{+}$ is a reflection order if and only if it coincides with one of the following orders when restricted to any dihedral reflection subgroup $(W,\{t_{1},t_{2}\})$:
@@
Let $\prec$ be a reflection order of the Coxeter system $(W,S)$ and Let $s\in S$. One can construct another reflection order, called the upper $s$-conjugate of $\prec$ and denoted by $\prec^{s}$, as follows: (1) if $\gamma_{1}\prec\gamma_{2}\prec\alpha_{s}$ and $\gamma_{1}\prec\gamma_{2}$, then $\gamma_{1}\prec^{s}\gamma_{2}$; (2) if $\alpha_{s}\prec\gamma_{1}\prec\gamma_{2}$, then $s(\gamma_{1})\prec^{s}s(\gamma_{2})$; (3) if $\gamma_{1}\prec\alpha_{s}\prec\gamma_{2}$, then $\gamma_{1}\prec^{s}\gamma_{2}$; (4) $\alpha_{s}$ is the maximum element under $\prec^{s}$. This notion is also introduced in [4]. See also Section 5.2 in [1].